# Resisting Quantum Key Distribution Attacks Using Quantum Machine Learning

### v1

@@ Abstract
The emergence of quantum computing poses significant risks to the security of modern communication networks as it breaks today’s public-key cryptographic algorithms. Quantum Key Distribution (QKD) offers a promising solution by harnessing the principles of quantum mechanics to establish secure keys. However, practical QKD implementations remain vulnerable to hardware imperfections and advanced attacks such as Photon Number Splitting and Trojan-Horse attacks. In this work, we investigate the potential of using quantum machine learning (QML) to detect popular QKD attacks. In particular, we propose a Hybrid Quantum Long Short-Term Memory (QLSTM) model to improve the detection of common QKD attacks. By combining quantum-enhanced learning with classical deep learning, the model captures complex temporal patterns in QKD data, improving detection accuracy. To evaluate the proposed model, we introduce a realistic QKD dataset simulating normal QKD operations along with seven attack scenarios, Intercept-and-Resend, Photon-Number Splitting (PNS), Trojan-Horse attacks Random Number Generator (RNG), Detector Blinding, Wavelength-dependent Trojan Horse, and Combined attacks. The dataset includes quantum security metrics such as Quantum Bit Error Rate (QBER), measurement entropy, signal and decoy loss rates, and time-based metrics, ensuring an accurate representation of real-world conditions. Our results demonstrate promising performance of the quantum machine learning approach compared to traditional classical machine learning models, highlighting the potential of hybrid techniques to enhance the security of future quantum communication networks. The proposed Hybrid QLSTM model achieved an accuracy of 93.7.0% after 50 training epochs, outperforming classical deep learning models such as LSTM, and CNN.
@@ I Introduction
The rise of quantum computing poses a significant challenge to the security of modern communication networks, particularly for encryption techniques based on mathematical problems. As quantum computers evolve, the infamous Shor’s algorithm will be able to break today’s asymmetric encryption techniques [1]. In response to these emerging threats, researchers have been working on Post-Quantum Cryptography (PQC) to develop encryption algorithms that can resist quantum attacks. PQC algorithms rely on complex mathematical structures, such as lattice-based, code-based, and multivariate polynomial cryptosystems, making them resistant to current quantum techniques. However, despite its promising security features, PQC remains reliant on mathematical problems, which means that its security is not provably unconditional. Given the rapid advancements in computational capabilities, even these new cryptographic approaches could face vulnerabilities in the future [2].
@@
An alternative and more radical direction is Quantum Key Distribution (QKD), which leverages the principles of quantum mechanics to generate keys that are theoretically secure against eavesdropping [1, 2]. Unlike classical key exchange methods, QKD benefits from the no-cloning theorem, which prevents copying quantum states. This property allows Alice (the sender) and Bob (the receiver) to detect any eavesdropping attempt, as such intrusion inevitably disturbs the transmitted quantum states [3]. This makes QKD attractive for sectors such as government, finance, and critical infrastructure, where extremely high levels of security are required [1].
@@
In recent years, QKD has progressed from theoretical studies to experimental prototypes and early-stage deployments. However, integrating QKD into existing communication infrastructures presents challenges related to scalability, network architecture, and resource optimization [4]. While global-scale QKD networks are being envisioned [5], real-world implementations remain vulnerable to practical threats. Hardware imperfections, channel noise, and side-channel attacks—such as Photon Number Splitting (PNS), Trojan Horse Attacks, and Channel Tampering—can all undermine the theoretical guarantees of QKD [5, 3]. These threats emphasize the importance of robust monitoring and intrusion detection strategies.
@@
Researchers have explored both conventional and data-driven approaches to securing QKD against such vulnerabilities. On the traditional side, non-ML methods focus on signal analysis, hardware countermeasures, and statistical detection. For example, spectral estimation techniques have been used to detect low-rate denial-of-service (LDoS) attacks in CV-QKD by identifying distinct low-frequency patterns in the power spectral density (PSD) [6]. Phase-sensitive amplifiers (PSA) combined with homodyne detection have been deployed to improve resilience against individual attacks [7], while plug-and-play architectures for Differential Phase Shift Measurement-Device-Independent QKD (DPS-MDI-QKD) help mitigate vulnerabilities in state preparation and measurement [8]. Network-layer defense strategies based on real-time traffic monitoring have also been proposed to detect denial-of-service attempts [9]. While these methods enhance QKD robustness, they are often limited by predefined assumptions, making them less adaptable to novel or evolving threats.
@@
This limitation has motivated a growing interest in Machine Learning (ML) and Deep Learning (DL) for QKD security. Unlike static threshold-based methods, ML models can adapt to dynamic patterns and detect both known and previously unseen attacks. In Continuous-Variable QKD (CV-QKD), Du and Huang [10] proposed multi-class detection using Binary Relevance Neural Networks (BR-NN) and Label Powerset Neural Networks (LP-NN), combined with one-class SVM for unknown attacks. Kish et al. [11] developed a decision tree–based framework for detecting and mitigating channel amplification attacks and related DoS scenarios. Mao et al. [12] used Artificial Neural Networks (ANN) to classify multiple attack types from optical parameter features such as quadrature values and local oscillator intensity. Luo et al. [13] introduced a semi-supervised GAN-based method that improved detection generalization to unseen threats. Table I summarizes these CV-QKD approaches.
@@
In Discrete-Variable QKD (DV-QKD), similar ML-based methods have been explored. Al et al. [14] applied ANN and LSTM models to detect Man-in-the-Middle attacks in IoT-integrated QKD networks. Tunc et al. [15] used LSTM and SVM classifiers to detect eavesdropping in quantum communication channels. Xu et al. [16] implemented Random Forest classifiers capable of detecting both device imperfections and eavesdropping in real-time. Table II summarizes these DV-QKD approaches.
@@
Despite these advances, existing ML approaches often rely on idealized simulations and limited attack scenarios. They tend to omit realistic conditions such as channel loss, phase noise, photon shot noise, depolarization, and detector imperfections, and often ignore critical quantum-level threats such as PNS, Intercept-and-Resend, Trojan Horse, RNG, Detector Blinding, and Wavelength-Dependent attacks. Furthermore, to the best of our knowledge, Quantum Machine Learning has yet to be applied to QKD intrusion detection.
To bridge this gap, we propose a hybrid intrusion detection system that integrates Quantum Long Short-Term Memory (QLSTM) into QKD monitoring. By combining quantum-enhanced learning with classical deep learning, the proposed model aims to improve detection accuracy and adaptability to evolving threats. We also introduce a new simulated dataset for decoy-state BB84 DV-QKD under eight scenarios: normal operation, Intercept-and-Resend, Trojan Horse, PNS, RNG, Wavelength-dependent Trojan Horse, Detector Blinding, and combined attacks.
@@ I-A Contributions
  • Comperhenisve QKD dataset: Given the lack of publicly available and comprehensive datasets, this work proposes a realistic QKD simulation using the PennyLane library. The dataset simulates decoy-state BB84 DVQKD under eight different scenarios: normal QKD operation, Intercept-and-Resend attack, Trojan Horse attack, PNS attack, RNG attack, Wavelength-dependent Trojan Horse attack, Detector Blinding attack, and combined attack. It also includes various quantum metrics such as Quantum Bit Error Rate (QBER), measurement entropy, signal and decoy loss rates, and time-based metrics, ensuring a dataset that accurately reflects real-world conditions.
  • Quantum Machine Learning for quantum intrusion detection on QKD systems: A hybrid intrusion detection model is proposed, integrating a QLSTM network with a classical LSTM. This approach leverages the advantages of both quantum computation and classical deep learning, offering a practical intrusion detection solution for future QKD networks.
@@ I-B Organization
The rest of this paper is organized as follows. Section II introduces the proposed Hybrid QLSTM model for detecting QKD attacks. Section III discusses the dataset generation process including the simulation algorithm that was used to generate the dataset. Section IV illustrates the evaluation setup and the used metrics. Section V presents and analyzes the experimental results, focusing on the performance of the proposed hybrid QLSTM model. Finally, Section LABEL:sec:Section_7 concludes this paper and provides suggestions for future work.
@@ II Model
This section presents the Hybrid QLSTM Model for QKD attack detection. Given the sequential nature of QKD transmissions, QLSTM is used to analyze quantum security metrics over time, improving attack detection accuracy compared to classical LSTM. The model consists of a QLSTM layer, a classical LSTM layer, and a fully connected layer. The QLSTM captures quantum-enhanced sequential features, the LSTM refines temporal dependencies, and the final layer maps features to logits for classification. The chapter also covers GPU-accelerated training, using CrossEntropy Loss, AdamW optimization, and a cosine annealing learning rate schedule. The model is trained for 50 epochs, with early stopping to prevent overfitting.
@@
In this work, we use QLSTM to analyze QKD security metrics over time and detect anomalies. QLSTM is an extension of the classical Long Short-Term Memory (LSTM) model, incorporating quantum principles, to enhance learning efficiency in sequential data tasks. Traditional LSTMs are widely used in time-series analysis, making them well-suited for detecting patterns in QKD-based attack scenarios. However, QLSTM leverages quantum mechanics to capture dependencies more effectively than classical models. As shown in [17], QLSTM outperformed classical LSTMs in detecting anomalies in Distributed Denial of Service (DDoS) attacks. Given the sequential nature of QKD transmissions, detecting anomalies in quantum metrics is crucial. Compared to classical LSTM, QLSTM converges faster and achieves higher accuracy in detecting attack patterns, making it a strong candidate for quantum security applications.
@@
The model consists of three main layers: a QLSTM layer, a classical LSTM layer, and a fully connected layer. Figure 1 depicts the high-level design of the proposed model. The QLSTM layer processes the sequential input data while incorporating quantum properties, such as entanglement and superposition, capturing complex features in the data. The output of the QLSTM layer is fed into a classical LSTM layer, which learns the temporal patterns from the quantum-enhanced sequences. Finally, a fully connected layer maps the learned features to the output space, producing logits, which are the scores that serve as the output of an ML model. These logits are passed to a softmax function for classification. Pseudocode 1 describes the model.
@@
II-A QLSTM Layer
This layer is implemented based on the open-source QLSTM implementation in [18]. However, for the purpose of this project, the original entanglement layer was changed from Basic Entanglement to Strongly Entanglement Layer. At each time step, the QLSTM receives the current input $x_{t}$ and the previous hidden and cell states, $h_{t}$ and $c_{t}$, respectively. It processes this information through variational quantum circuits (VQCs), representing the Forget Gate, Input Gate, Candidate Cell State, and Output Gate. Specifically, the gates are computed as follows:
$i_{t}$ $=\sigma(VQC_{2}(v_{t}))\quad f_{t}=\sigma(VQC_{1}(v_{t})),$ (1)
$\tilde{C}_{t}$ $=\tanh(VQC_{3}(v_{t})),\quad o_{t}=\sigma(VQC_{4}(v_{t})),$ (2)
@@ II-A QLSTM Layer
This information is processed through variational quantum circuits (VQCs), Forget Gate, Input Gate, Update Gate, and Output Gate. The outputs of the VQCs are passed through classical nonlinearities ($\sigma$, $\tanh$) before combining to produce the updated hidden and cell states. These updated states are then propagated to the next time step. The architecture of the QLSTM model used in this work is visually illustrated in Figure 2 (a).
II-B Classical LSTM Layer
This layer consists of 32 hidden states and operates in three main steps:
@@ II-B Classical LSTM Layer
At each time step $t$, the LSTM layer updates its internal memory using the input $x_{t}$ and the previous hidden and cell states ($h_{t-1}$, $c_{t-1}$). As illustrated in Figure 2 (b), this process involves three gates, the forget gate $f_{t}$, input gate $i_{t}$, and output gate $o_{t}$, along with a candidate cell state $\tilde{c}_{t}$. The gate activations are computed as follows:
@@
This ensures that the output at each step is a filtered representation of the current cell state. After processing the entire sequence, the final hidden state $h_{t}$ is extracted and passed into the next stage for classification.
II-C Fully Connected Layer
The fully connected layer maps output from the classical LSTM into the final output space, which consists of 4 logits, each corresponding to a possible class (Normal, Intercept-and-Resend Attack, PNS Attack, Trojan-Horse Attack). These logits are later transformed to class probabilities using a softmax function during evaluation.

### v2

@@ Abstract
The emergence of quantum computing poses significant risks to the security of modern communication networks as it breaks today’s public-key cryptographic algorithms. Quantum Key Distribution (QKD) offers a promising solution by harnessing the principles of quantum mechanics to establish secure keys. However, practical QKD implementations remain vulnerable to hardware imperfections and advanced attacks such as Photon Number Splitting and Trojan-Horse attacks. In this work, we investigate the potential of quantum machine learning (QML) to detect popular QKD attacks. In particular, we propose a Hybrid Quantum Long Short-Term Memory (QLSTM) model to improve the detection of common QKD attacks. By combining quantum-enhanced learning with classical deep learning, the model captures complex temporal patterns in QKD data, improving detection accuracy. To evaluate the proposed model, we introduce a realistic QKD dataset that simulates typical QKD operations, along with seven attack scenarios: Intercept-and-Resend, Photon-Number Splitting (PNS), Trojan-Horse attacks on the Random Number Generator (RNG), Detector Blinding, Wavelength-dependent Trojan-Horse, and Combined attacks. The dataset includes quantum security metrics such as Quantum Bit Error Rate (QBER), measurement entropy, signal and decoy loss rates, and time-based metrics, ensuring an accurate representation of real-world conditions. Our results demonstrate the promising performance of the quantum machine learning approach compared to traditional classical machine learning models, highlighting the potential of hybrid techniques to enhance the security of future quantum communication networks. The proposed Hybrid QLSTM model achieved an accuracy of 94.7.0% after 50 training epochs, outperforming classical deep learning models such as LSTM, CNN, ANN, Random Forest, and RNN. However, our evaluation is conducted on a semi-realistic, simulation-generated decoy-state BB84 dataset with simplified models of attacks and device imperfections, so the reported performance should be interpreted as a proof-of-concept rather than a final assessment on field-deployed QKD systems.
@@ I Introduction
The rise of quantum computing poses a significant challenge to the security of modern communication networks, particularly for encryption techniques based on mathematical problems. As quantum computers evolve, the infamous Shor’s algorithm will be able to break current asymmetric encryption techniques [8]. In response to these emerging threats, researchers have been developing Post-Quantum Cryptography (PQC) to create encryption algorithms that resist quantum attacks. PQC algorithms rely on complex mathematical structures, such as lattice-based, code-based, and multivariate polynomial cryptosystems, making them resistant to current quantum techniques. However, despite its promising security features, PQC remains reliant on mathematical problems, so its security is not provably unconditional. Given the rapid advances in computational capabilities, even these new cryptographic approaches could face vulnerabilities in the future [18].
@@
An alternative and more radical direction is the Quantum Key Distribution (QKD), which uses principles of quantum mechanics to generate keys that are theoretically secure against eavesdropping [8, 18]. Unlike classical key exchange methods, QKD benefits from the no-cloning theorem, which prevents the cloning of quantum states. This property allows Alice (the sender) and Bob (the receiver) to detect any eavesdropping attempt, as such an intrusion inevitably disturbs the transmitted quantum states [10]. This makes QKD attractive for sectors such as government, finance, and critical infrastructure, where extremely high levels of security are required [8].
@@
In recent years, QKD has progressed from theoretical studies to experimental prototypes and early-stage deployments. However, integrating QKD into existing communication infrastructures presents challenges related to scalability, network architecture, and resource optimization [3]. Although global-scale QKD networks are being envisioned [15], real-world implementations remain vulnerable to practical threats. Hardware imperfections, channel noise, and side-channel attacks—such as Photon Number Splitting (PNS), Trojan Horse Attacks, and Channel Tampering—can undermine the theoretical guarantees of QKD [15, 10]. These threats underscore the importance of robust monitoring and intrusion-detection strategies.
@@
Researchers have explored both conventional and data-driven approaches to securing QKD against such vulnerabilities. On the traditional side, non-ML methods focus on signal analysis, hardware countermeasures, and statistical detection. For example, spectral estimation techniques have been used to detect low-rate denial-of-service (LDoS) attacks in CV-QKD by identifying distinct low-frequency patterns in the power spectral density (PSD) [4]. Phase-sensitive amplifiers (PSA) combined with homodyne detection have been deployed to improve resilience against individual attacks [2], while plug-and-play architectures for Differential Phase Shift Measurement-Device-Independent QKD (DPS-MDI-QKD) help mitigate vulnerabilities in state preparation and measurement [26]. Network-layer defense strategies based on real-time traffic monitoring have also been proposed to detect denial-of-service attempts [19]. Although these methods improve QKD robustness, they are often constrained by predefined assumptions, limiting their adaptability to novel or evolving threats.
@@
This limitation has motivated a growing interest in Machine Learning (ML) and Deep Learning (DL) for QKD security. Unlike static threshold-based methods, ML models can adapt to dynamic patterns and detect known and previously unseen attacks. In Continuous-Variable QKD (CV-QKD), Du and Huang [5] proposed multi-class detection using Binary Relevance Neural Networks (BR-NN) and Label Powerset Neural Networks (LP-NN), combined with one-class SVM for unknown attacks. Kish et al. [16] developed a decision tree–based framework to detect and mitigate channel amplification attacks and related DoS scenarios. Mao et al. [24] used Artificial Neural Networks (ANN) to classify multiple attack types from optical parameter features such as quadrature values and local oscillator intensity. Luo et al. [21] introduced a semi-supervised GAN-based method that improved the detection generalization of unseen threats. Table I summarizes these CV-QKD approaches.
@@
In Discrete-Variable QKD (DV-QKD), similar ML-based methods have been explored. Al et al. [1] applied ANN and LSTM models to detect Man-in-the-Middle attacks in QKD networks integrated with IoT. Tunc et al. [30] used LSTM and SVM classifiers to detect eavesdropping in quantum communication channels. Xu et al. [31] implemented Random Forest classifiers capable of detecting both device imperfections and eavesdropping in real-time. Table II summarizes these DV-QKD approaches.
@@
Despite these advances, existing ML approaches often rely on idealized simulations and limited attack scenarios. They tend to omit realistic conditions such as channel loss, phase noise, photon shot noise, depolarization, and detector imperfections, and often ignore critical quantum-level threats such as PNS, Intercept-and-Resend, Trojan Horse, RNG, detector blinding, and Wavelength-Dependent attacks. Furthermore, to the best of our knowledge, Quantum Machine Learning has not yet been applied to QKD intrusion detection.
To bridge this gap, we propose a hybrid intrusion detection system that integrates Quantum Long Short-Term Memory (QLSTM) into QKD monitoring. By combining quantum-enhanced learning with classical deep learning, the proposed model aims to improve detection accuracy and adaptability to evolving threats. We also introduce a new simulated dataset for the decoy-state BB84 DV-QKD under eight scenarios: normal operation, Intercept-and-Resend, Trojan Horse, PNS, RNG, Wavelength-dependent Trojan Horse, Detector Blinding, and combined attacks.
@@ I-A Contributions
  • Comprehensive QKD dataset: Given the lack of publicly available and comprehensive datasets, this work proposes a realistic QKD simulation using the PennyLane library. The dataset simulates decoy-state BB84 DVQKD under eight different scenarios: normal QKD operation, Intercept-and-Resend attack, Trojan Horse attack, PNS attack, RNG attack, Wavelength-dependent Trojan Horse attack, Detector Blinding attack, and combined attack. It also includes various quantum metrics such as Quantum Bit Error Rate (QBER), measurement entropy, signal and decoy loss rates, and time-based metrics, ensuring a dataset that accurately reflects real-world conditions.
  • Quantum Machine Learning for quantum intrusion detection on QKD systems: A hybrid intrusion detection model is proposed, integrating a QLSTM network with a classical LSTM. This approach leverages the strengths of both quantum computation and classical deep learning, offering a practical intrusion detection solution for future QKD networks.
@@ I-B Organization
The remainder of this paper is organized as follows. Section II introduces the proposed Hybrid QLSTM model to detect QKD attacks. Section III discusses the dataset generation process, including the simulation algorithm that was used to generate the dataset. Section IV illustrates the evaluation setup and the metrics used. Section V presents and analyzes the experimental results, focusing on the performance of the proposed hybrid QLSTM model. Finally, Section LABEL:sec:Section_7 concludes this paper and provides suggestions for future work.
@@ II Model
This section presents the Hybrid QLSTM Model for QKD attack detection. Given the sequential nature of QKD transmissions, QLSTM is used to analyze quantum security metrics over time, thereby improving attack detection accuracy compared to classical LSTM. The model consists of a QLSTM layer, a classical LSTM layer, and a fully connected layer. The QLSTM captures quantum-enhanced sequential features, the LSTM refines temporal dependencies, and the final layer maps features to classification logits. The section also covers GPU-accelerated training using CrossEntropyLoss, AdamW optimization, and a cosine annealing learning rate schedule. The model is trained for 50 epochs, with early stopping to prevent overfitting.
@@
In this work, we use QLSTM to analyze QKD security metrics over time and detect anomalies. QLSTM is an extension of the classical Long Short-Term Memory (LSTM) model that incorporates quantum principles to enhance learning efficiency in sequential data tasks. Traditional LSTMs are widely used in time-series analysis, making them well-suited to detect patterns in QKD-based attack scenarios. However, QLSTM leverages quantum mechanics to capture dependencies more effectively than classical models. As shown in [29], QLSTMs outperformed classical LSTMs in detecting anomalies in Distributed Denial of Service (DDoS) attacks. Given the sequential nature of QKD transmissions, detecting anomalies in quantum metrics is crucial. Compared to classical LSTMs, QLSTMs converge faster and achieve higher accuracy in detecting attack patterns, making them a strong candidate for quantum security applications.
II-A Foundations of LSTM and QLSTM
Traditional ML models, such as LSTM, are designed to process sequential data by learning temporal dependencies through gated recurrent units. LSTM networks have been successfully applied in various anomaly-detection and time-series tasks; however, their representational capacity is bounded by classical computational limits. QML extends these methods by embedding data in quantum states and processing them through variational quantum circuits (VQCs). These circuits exploit quantum mechanical principles, such as superposition, entanglement, and interference, to represent and manipulate complex correlations in fewer dimensions. This provides a potential advantage in optimization efficiency. The QLSTM model combines these properties and replaces or augments the classical LSTM gates with variational quantum circuits, enabling the model to learn temporal relationships within a quantum-enhanced feature space. This hybrid structure allows quantum layers to capture complex dependencies in the input sequence. In contrast, classical layers handle large-scale temporal aggregation, thereby improving convergence and generalization in sequential learning tasks.
Quantum circuits operate within a high-dimensional Hilbert space, where qubit superposition and entanglement allow data to be represented in exponentially richer feature spaces compared to classical vector encoding [14]. This enables more expressive temporal feature extraction with fewer parameters, leading to improved optimization landscapes and enabling faster gradient-based training. Furthermore, many studies have shown that QLSTM models exhibit accelerated convergence and enhanced prediction accuracy in various time-series domains, including renewable energy forecasting [14] and cybersecurity [29]. In these studies, QLSTM consistently outperformed its classical LSTM counterpart, achieving lower validation loss within early epochs while maintaining superior generalization on unseen samples. These results empirically confirm that quantum recurrent layers can mitigate vanishing-gradient effects and capture complex nonlinear dependencies more effectively than purely classical architectures.
II-B Model Architecture
@@
The model consists of three main layers: a QLSTM layer, a classical LSTM layer, and a fully connected layer. Figure 1 shows the high-level design of the proposed model. The QLSTM layer processes the sequential input data while incorporating quantum properties, such as entanglement and superposition, capturing complex features in the data. The output of the QLSTM layer is fed into a classical LSTM layer, which learns the temporal patterns from the quantum-enhanced sequences. Finally, a fully connected layer maps the learned features to the output space, producing logits, which are the scores that serve as the output of an ML model. These logits are passed to a softmax function for classification. Pseudocode 1 describes the model.
@@
II-B1 QLSTM Layer
This layer is implemented based on the open-source QLSTM implementation in [27]. However, for the purpose of this project, the original entanglement layer was changed from Basic Entanglement to Strongly Entanglement Layer. At each time step, the QLSTM receives the current input $x_{t}$ and the previous hidden and cell states, $h_{t}$ and $c_{t}$, respectively. It processes this information using VQCs that represent the Forget Gate, the Input Gate, the Candidate Cell State, and the Output Gate. Each VQC operates on qubits that use superposition and entanglement to encode information in a high-dimensional Hilbert space. This enables the gates to represent and learn complex feature relationships with fewer parameters than classical layers, improving training efficiency and representation power. Specifically, the gates are computed as follows:
$i_{t}$ $=\sigma(VQC_{2}(v_{t}))\qquad f_{t}=\sigma(VQC_{1}(v_{t})),$ (1)
$\tilde{C}_{t}$ $=\tanh(VQC_{3}(v_{t})),\qquad o_{t}=\sigma(VQC_{4}(v_{t})),$ (2)
@@ II-B1 QLSTM Layer
The outputs of the VQCs are passed through classical nonlinearities ($\sigma$, $\tanh$) before combining to produce the updated hidden and cell states. These updated states are then propagated to the next time step. The architecture of the QLSTM model used in this work is visually illustrated in Figure 2 (a).
II-B2 Classical LSTM Layer
The LSTM captures long-term and short-term dependencies in sequential data through its cell and hidden states. It regulates information flow using input, forget, and output gates, which decide what to keep or discard from previous time steps. In this model, the LSTM refines the temporal features extracted from the QLSTM, ensuring stable learning and smoother transitions before classification. In this proposed model, the LSTM layer consists of 32 hidden states and operates in three main steps:
@@ II-B2 Classical LSTM Layer
At each time step $t$, the LSTM layer updates its internal memory using input $x_{t}$ and the previous hidden and cell states ($h_{t-1}$, $c_{t-1}$). As illustrated in Figure 2 (b), this process involves three gates, the forget gate $f_{t}$, the input gate $i_{t}$, and the output gate $o_{t}$, along with a candidate cell state $\tilde{c}_{t}$. Gate activations are computed as follows:
@@
This ensures that the output at each step is a filtered representation of the current cell state. After processing the entire sequence, the final hidden state $h_{t}$ is extracted and passed to the next stage for classification. This process allows the model to capture long-term temporal dependencies while mitigating vanishing gradients.
II-B3 Fully Connected Layer
The fully connected layer maps the classical LSTM’s output to the final output space, which consists of 4 logits, each corresponding to a possible class (Normal, Intercept-and-Resend Attack, PNS Attack, Trojan-Horse Attack). These logits are later transformed to class probabilities using a softmax function during evaluation.